# https://vision.ist.i.kyoto-u.ac.jp/

### v2

@@ 4.2 Silhouette Projector
In the process of creating a scaled camera height supervision, we need to know the per-frame scale factor. For this, we can consider leveraging an object of known size, i.e., we can use a prior on the object size (i.e., dimensions). To achieve this, we will need to know the 3D region which defines object dimensions. This is a non-trivial task as objects are often occluded or cropped out of images and they are in various poses.
We introduce Silhouette Projector, which measures unscaled object heights from the estimated depth map in a simple but robust manner. The scale factor can then be computed by comparing them with the height priors. The main idea of Silhouette Projector is based on two facts: 1) the height of the object silhouette projected on the plane perpendicular to the ground plane does not change regardless of the object’s pose, and 2) we can compute the silhouette height by measuring the distance from the top of the silhouette to the ground plane even if they are partly occluded or truncated, as long as the top is visible and the object is on the ground plane.
We first reconstruct an object point cloud with the estimated depth map through perspective reprojection. The object region $M_{\mathrm{obj}}$ in the image can be obtained with an off-the-shelf instance segmentation model [23] before training. We then orthographically project these 3D points onto a plane $P^{\perp}$ perpendicular to the ground plane, i.e., an arbitrary plane parallel to the road normal vector $\mathbf{\tilde{n}}$. We compute $\mathbf{\tilde{n}}$ as the median of the normal vectors $\hat{\mathbf{n}}(p_{i})$ (see Eq. 6) in the road region $M_{\mathrm{r}}$. We define the farthest distance from the 2D point on $P^{\perp}$ to the ground plane as the silhouette height. Keep in mind that this value is unscaled because we derived this from the estimated unscaled depth map. The ratio between a height prior and its silhouette height corresponds to the scale factor of interest. When there are multiple objects in an image, we use the median of the scale factors each corresponding to each object as the scale factor of the frame for computational simplicity and robustness.
@@
Some objects cannot be handled well with Silhouette Projector for various reasons including occluded upper parts, not on the ground plane, inaccurate segmented region, and misclassified object category. To filter these as outliers when computing the scale factors, we introduce an algorithm to detect them based on geometric plausibility. We first calculate the horizon $\mathbf{l}$ with the road normal vector $\mathbf{\tilde{n}}$ and the camera intrinsics $\mathbf{K}$ [17]:
@@
Based on the approximations in [20], we derive the approximate metric-scale object height $\hat{H}_{\mathrm{obj}}\approx\frac{h_{\mathrm{obj}}}{h_{\mathrm{cam}}}H^{*\tau-1}_{\mathrm{cam}}$ where $h_{\mathrm{cam}}$ is the farthest distance from the pixel in the object instance mask to $\mathbf{l}$, $h_{\mathrm{obj}}$ is the height of the 2D bounding box, and $H^{*\tau-1}_{\mathrm{cam}}$ is the pseudo camera height at the $\tau$-th epoch derived in Eq. 9. By comparing the relative gap between $\hat{H}_{\mathrm{obj}}$ and the object height prior $H_{\mathrm{obj}}$, we determine the outliers as
@@ 4.3 Training Losses
We experimentally observe that estimated object depths, especially those distant from the camera, often erroneously lean towards the road surface because of the inaccurate supervisory signal from the reconstruction loss in low-texture regions. This tendency makes silhouette heights inaccurate. To mitigate this issue, we introduce an auxiliary rough geometric loss $\mathcal{L}_{\mathrm{aux}}$, which enforces object point clouds to be within a close region. For each inlier object $k\,(k=1\dots K)$, we define $\mathcal{L}_{\mathrm{aux}}$ as the gap between the estimated depth map $D$ in the object region $M_{\mathrm{obj}}$ and the approximate object depth $D_{\mathrm{aprx}}=\frac{H_{\mathrm{obj}}}{h_{\mathrm{obj}}}f_{y}$ where $f_{y}$ is the focal length:
@@
The $\mathcal{L}_{\mathrm{aux}}$ is potentially inaccurate since $D_{\mathrm{aprx}}$ assumes that each object can be represented by a plane parallel to the image plane, and we observe that it degrades the depth accuracy in the late stage of training. On the other hand, in the early stage, an estimated depth is unreliable, so heavily relying on the camera height loss $\mathcal{L}_{\mathrm{cam}}$ makes the training unstable. For this, we gradually decrease the weight of $\mathcal{L}_{\mathrm{aux}}$ and increase the one for $\mathcal{L}_{\mathrm{cam}}$ as the training proceeds in a logarithmic way. To ensure training stability, we stop changing these weights after the $\tau_{\mathrm{mid}}$-th epoch. This balancing can be expressed as:
@@ 5 LSP
Our StableCamH extracts metric scale from a prior on object heights through the camera height loss $\mathcal{L}_{\mathrm{cam}}$ and the auxiliary rough geometric loss $\mathcal{L}_{\mathrm{aux}}$. To ensure that StableCamH learns the scale regardless of the particular dataset used for training, a universal prior of object heights, independent of the environment, becomes essential. Cars are objects whose dimensions are explicitly and accurately known. Better yet, they are abundantly located on the road, the key semantic region whose metric depth is essential for many downstream tasks. Some methods for other tasks have modeled the size prior as a probability distribution [20, 60]. The gap between its mean and the true expectation for each environment, however, is directly reflected in the uncertainty of the scale in the supervision. Wang et al. [42] also report that there is a gap in car dimensions among datasets. These facts suggest that modeling the object size prior as a single probability distribution is problematic, especially for man-made objects like cars.
@@
For training, we use a dataset consisting of a large number of vehicle images and their information such as 3D dimensions. This dataset can be easily collected with web scraping and does not require manual annotations. To enhance the robustness to various input images, in addition to common color jittering and blur augmentation, we perform augmentations to simulate occlusion and truncation. For occlusion augmentation, we mask out random regions with the same shape as a vehicle or with a beam of random width. For truncation augmentation, we translate the vehicle mask in a random direction. We visualize these in supplementary material. The abundance of training images and diverse augmentations enable LSP to estimate vehicle dimensions robustly across various datasets.
Although StableCamH requires only a car height prior, we construct LSP to estimate the width and length as well as height. This enhances it to grasp an overall car size and improves the prediction accuracy of car heights.
@@ 6 Experiments
6.1 StableCamH
We evaluate the accuracy of depth estimators trained on our framework and demonstrate the effectiveness of StableCamH as a novel unsupervised metric-scale training method through depth prediction accuracy, its applicability to any network architecture, and its ability to leverage any dataset for training.
@@ 6.1 StableCamH
We implement StableCamH with PyTorch [30]. We employ Monodepth2 [15] with ResNet50 encoder pre-trained on ImageNet [32] as a default depth estimator which is one of the most basic architectures in self-supervised MDE. For all experiments unless mentioned otherwise, we train the model for 50 epochs with a batch size of 8, and set $\tau_{\mathrm{mid}}$ to 20. We use Adam optimizer [25] with an initial learning rate of $5.0\times 10^{-5}$ decayed by half every 15 epochs. During training, we perform the following augmentations with 50% chance in random order: horizontal flips, brightness adjustment ($\pm$0.2), saturation adjustment ($\pm$0.2), contrast adjustment ($\pm$0.2), and hue jitter ($\pm$0.1). For quantitative evaluation, we follow [9] and compute the seven standard metrics: AbsRel, SqRel, RMSE, $\mathrm{RMSE_{log}}$, $\delta<1.25$, $\delta<1.25^{2}$, and $\delta<1.25^{3}$.
Table 1: Quantitative results on the KITTI Eigen test split [8] (640$\times$192 resolution). In the Supervision column, V, G, and CamH represents velocity, gravity, and camera height, respectively. The Scaling column represents whether to perform the median scaling [59]. We denote the best results in bold and the second best ones with an underline for each block. Note that we train VADepth [48] from scratch using ground-truth camera height (1.65m) provided by KITTI unlike its original paper. StabeCamH outperforms weakly-supervised methods regardless of the network architectures and without the need for ground truth supervision. In the rescaled results, every model trained with StableCamH achieves better results than its original one.
@@
[…]
@@
Figure 2: Qualitative comparison on KITTI. In error maps, the larger depth errors are represented in red, smaller ones in blue. The model trained with our StableCamH predicts more accurate depth map compared to the weakly-supervised methods.
@@
Our StableCamH is independent of network architectures so we experiment on several models with StableCamH as well as Monodepth2: Lite-Mono [55] with ResNet18, HR-Depth [27] with ResNet18, and VADepth [48]. We train these models on the KITTI raw dataset [12], which contains 39,810 images for training and 4,424 for validation, and evaluate it with the KITTI Eigen test split [8] (697 images). For training and evaluation, all images are resized to 640$\times$192.
We show the qualitative results in Fig. 2. The quantitative results shown in Tab. 1 demonstrate that StableCamH enables the various architectures to learn metric scale properly. Even when employing the simplest network architecture, i.e., Monodepth2, it outperforms weakly-supervised methods that rely on ground-truth scale labels for training which our method does not need. Similar to VADepth [48], StableCamH uses the camera height invariance as supervision though VADepth requires the ground-truth and employs a complex network architecture. Nevertheless, StableCamH with Monodepth2 (row 6 in Tab. 1) surpasses it. This indicates that measuring camera height precisely is difficult and using unreliable ground-truth camera height causes low accuracy in the scale of estimated depth maps.
In the results showing the accuracy of the estimated depth map scaled with ground-truth scale factors via median scaling [59], it is evident that every model trained with StableCamH attains superior results to the one trained without StableCamH. This shows that our training framework not only enables the models to learn metric scale but also enhances the geometric accuracy. This improvement is thanks to the camera height loss $\mathcal{L}_{\mathrm{cam}}$ and $\mathcal{L}_{\mathrm{aux}}$, which makes the model respect road regions and object regions, respectively.
@@
To evaluate the ability of StableCamH for taking advantage of any dataset for training, we train and evaluate it on the Cityscapes dataset [6]. We follow [59, 49, 51, 45] and collect 69,731 images for training and evaluate on the 1,525 test images. In this dataset, there are multiple camera intrinsics setups. Since this strongly affects the metric depth estimation [50], we first align the focal lengths of all images to 587.5 by resizing them, then crop them to 512$\times$192 resolution. For evaluation, we follow [2, 45] and further crop the estimated image into 416$\times$128. In order to align the number of iterations on KITTI, we train for 29 epochs with the same learning rate decreased by half every 8 epochs, and $\tau_{\mathrm{mid}}$ is set to 12.
Tab. 3 shows the quantitative results and StableCamH outperforms weakly-supervised methods by wide margins. These methods assume that highly accurate sensor data is available. Therefore, they fail to train scale-aware MDE models on datasets consisting of unreliable sensor data such as Cityscapes. On the other hand, StableCamH does not require any sensor measurement other than RGB video, so it can robustly leverage any driving videos for training.
@@
To demonstrate that StableCamH can train a model on a mixed set of datasets to attain strong generalizability, we train with 235,341 images collected from the Argoverse2 [46], Lyft [21], A2D2 [13], and DDAD [16] dataset and evaluate it on the KITTI Eigen test split [8]. Both images for training and evaluation are resized and cropped to 832$\times$512 resolution. The training images are the sets of video sequences, with each sequence being captured with its own unique camera setting. We individually optimize each camera height for each sequence during training. We detail the dataset composition and training settings in supplementary material.
As shown in Tab. 3, the model trained with StableCamH on the large-scale mixed dataset achieves comparable accuracy to the one trained on KITTI despite the domain gap and attains high generalizability. The model trained on both the mixed dataset and KITTI outperforms the model trained only on KITTI. This suggests that training on various images enhances accuracy which is the strength of fully self-supervised learning methods.
Table 2: Quantitative results on Cityscapes [6]. We train all the models from scratch with the same supervisions as the Supervision column in Tab. 1. StableCamH significantly outperforms weakly-supervised methods.
Table 3: StableCamH can be trained on mixed datasets (Argoverse2, Lyft, A2D2, and DDAD) with various camera settings, and attains high generalizability. The results are evaluated on the KITTI Eigen split [8] (832$\times$512 resolution).
@@
StableCamH 0.125 1.288 6.359 0.858
@@
Table 3: StableCamH can be trained on mixed datasets (Argoverse2, Lyft, A2D2, and DDAD) with various camera settings, and attains high generalizability. The results are evaluated on the KITTI Eigen split [8] (832$\times$512 resolution).
@@
To study the effect of the proposed components in StableCamH, we conduct ablation studies on the training losses, balancing the loss weights, camera height optimization, and outlier filtering. The results are shown in Tab. 5. Thanks to the accuracy of LSP and the robustness of Silhouette Projector, the camera height loss makes a greater contribution to the depth accuracy than the auxiliary rough geometric loss. By introducing the camera height optimization, the model can learn the scale with more accurate and consistent supervision than leveraging the size prior independently for each frame (rows 1, 2, and 6 in Tab. 5), which leads to improved accuracy. Though balancing the loss weights (Eq. 14) with either $\mathcal{L}_{\mathrm{aux}}$ or $\mathcal{L}_{\mathrm{cam}}$ degrades the results due to the insufficient strength of the supervisory signals for metric scale, it contributes to improving accuracy when both $\mathcal{L}_{\mathrm{aux}}$ and $\mathcal{L}_{\mathrm{cam}}$ are employed.
Our proposed method computing the camera height pseudo-supervision from estimated depth maps with Silhouette Projector and LSP can also be used with a pre-trained MDE model in an offline manner. We first train the original Monodepth2 [15] without StableCamH, then predict depth maps of training images. With Silhouette Projector and LSP, we obtain per-frame scaled camera height and optimize it across frames with the median operation. We fix this value and use it as the camera height supervision for an untrained Monodepth2 on StableCamH. We examine the difference between the proposed camera height optimized online and the fixed one derived offline when we use them as the pseudo-supervision. Note that we also fix $\lambda_{\mathrm{cam}}$ because the pseudo-label can be considered sufficiently accurate from the early stage of training. We also run another experiment in which we fix camera height supervision derived offline for the first 20 epochs, which is the same number of training epochs as the original Monodepth2, then unfix the camera height pseudo-label to fine-tune it via online optimization. As shown in Tab. 5, the model supervised with the fixed camera height computed offline achieves slightly better results than the one supervised with the camera height optimized online. This result indicates that our strategy of deriving scaled camera height is highly accurate and fixing it stabilizes the training. The model trained with the fine-tuned camera height achieves the highest accuracy, which means our StableCamH can attain better accuracy with additional pre-training costs.

### v3

@@ 4.2 Silhouette Projector
In the process of creating a scaled camera height supervision, we need to know the per-frame scale factor. For this, we can consider leveraging an object of known size. This, however, would necessitate extraction of the exact 3D object region which is non-trivial as objects are often occluded or cropped and oriented in various poses.
We introduce Silhouette Projector to robustly estimate this per-frame scale factor. Silhouette Projector leverages two simple facts of road scenes. First, the height of the object silhouette projected on the plane perpendicular to the ground plane does not change regardless of the object pose. Second, we can compute the silhouette height by measuring the distance from the top of the silhouette to the ground plane even if they are partly occluded or truncated, as long as the top is visible and the object is on the ground plane.
We first reconstruct an object point cloud with the estimated depth map through perspective reprojection. The object region $M_{\mathrm{obj}}$ in the image can be obtained with an off-the-shelf instance segmentation model [26] before training. We then orthographically project these 3D points onto a plane $P^{\perp}$ perpendicular to the ground plane, i.e., an arbitrary plane parallel to the road normal vector $\mathbf{\tilde{n}}$. We compute $\mathbf{\tilde{n}}$ as the median of the normal vectors $\hat{\mathbf{n}}(p_{i})$ (see Eq. 6) in the road region $M_{\mathrm{r}}$. We define the farthest distance from the 2D point on $P^{\perp}$ to the ground plane as the silhouette height. This value is unscaled because we derived this from the estimated unscaled depth map. The ratio between a height prior and its silhouette height corresponds to the scale factor of interest. When there are multiple objects in an image, we use the median of the scale factors each corresponding to each object as the scale factor of the frame for computational simplicity and robustness.
@@
Some objects cannot be handled well with Silhouette Projector for various reasons including occluded upper parts, not on the ground plane, inaccurate segmented region, and misclassified object category. We introduce an algorithm to detect these as outliers based on geometric plausibility. We first calculate the horizon $\mathbf{l}$ with the road normal vector $\mathbf{\tilde{n}}$ and the camera intrinsics $\mathbf{K}$ [20]
@@
Based on the approximations in [23], we derive the approximate metric-scale object height $\hat{H}_{\mathrm{obj}}\approx\frac{h_{\mathrm{obj}}}{h_{\mathrm{cam}}}H^{*\tau-1}_{\mathrm{cam}}$, where $h_{\mathrm{cam}}$ is the farthest distance from the pixel in the object instance mask to $\mathbf{l}$, $h_{\mathrm{obj}}$ is the farthest distance between two points on the object mask on which tangent lines parallel to $\mathbf{l}$ lie, and $H^{*\tau-1}_{\mathrm{cam}}$ is the pseudo camera height at the $\tau$-th epoch derived in Eq. 9. By comparing the relative gap between $\hat{H}_{\mathrm{obj}}$ and the object height prior $H_{\mathrm{obj}}$, we determine the outliers as
@@ 4.3 Training Losses
We experimentally observe that estimated object depths, especially those distant from the camera, often erroneously lean towards the road surface because of the inaccurate supervisory signal from the reconstruction loss in low-texture regions. This tendency makes silhouette heights inaccurate. To mitigate this issue, we introduce an auxiliary rough geometric loss $\mathcal{L}_{\mathrm{aux}}$, which enforces object point clouds to be within a close region. For each inlier object $k\,(k=1\dots K)$, we define $\mathcal{L}_{\mathrm{aux}}$ as the gap between the estimated depth map $D$ in the object region $M_{\mathrm{obj}}$ and the approximate object depth $D_{\mathrm{aprx}}=\frac{H_{\mathrm{obj}}}{h_{\mathrm{obj}}}f_{y}$ where $f_{y}$ is the focal length
@@
The $\mathcal{L}_{\mathrm{aux}}$ is potentially inaccurate since $D_{\mathrm{aprx}}$ assumes that each object can be represented by a plane parallel to the image plane, and we observe that it degrades the depth accuracy in the late stage of training. On the other hand, in the early stage, estimated depth is unreliable, so heavily relying on the camera height loss $\mathcal{L}_{\mathrm{cam}}$ makes the training unstable. For this, we gradually decrease the weight of $\mathcal{L}_{\mathrm{aux}}$ and increase the one for $\mathcal{L}_{\mathrm{cam}}$ as the training proceeds in a logarithmic way. To ensure training stability, we stop changing these weights after the $\tau_{\mathrm{mid}}$-th epoch
@@ 5 LSP
Our FUMET extracts metric scale from a prior on object heights through the camera height loss $\mathcal{L}_{\mathrm{cam}}$ and the auxiliary rough geometric loss $\mathcal{L}_{\mathrm{aux}}$. To ensure that FUMET learns the scale regardless of the particular dataset used for training, a universal prior of object heights, independent of the environment, becomes essential. Cars are objects whose dimensions are explicitly and accurately known. Better yet, they are abundantly located on the road, the key semantic region whose metric depth is essential for many downstream tasks. Some methods for other tasks have modeled the size prior as a probability distribution [23, 64]. The gap between its mean and the true expectation for each environment, however, is directly reflected in the uncertainty of the scale in the supervision. Wang et al. [46] also report that there is a gap in car dimensions among datasets. These facts suggest that modeling the object size prior as a single probability distribution is problematic, especially for man-made objects like cars.
@@
For training, we use a dataset consisting of a large number of vehicle images and their information such as 3D dimensions. This dataset can be easily collected with web scraping and does not require manual annotations. To enhance the robustness to various input images, in addition to common color jittering and blur augmentation, we perform augmentations to simulate occlusion and truncation. For occlusion augmentation, we mask out random regions with the same shape as a vehicle or with a beam of random width. For truncation augmentation, we translate the vehicle mask in a random direction. We visualize these in the supplementary material. The abundance of training images and diverse augmentations enable LSP to estimate vehicle dimensions robustly across various datasets.
Although FUMET requires only a car height prior, we construct LSP to estimate the width and length as well as height. This enhances it to grasp an overall car size and improves the prediction accuracy of car heights.
@@ 6 Experiments
6.1 FUMET
We evaluate the accuracy of depth estimators trained on our framework and demonstrate the effectiveness of FUMET as a novel unsupervised metric-scale training method through depth prediction accuracy, its applicability to any network architecture, and its ability to leverage any dataset for training.
@@ 6.1 FUMET
We implement FUMET with PyTorch [33]. We employ Monodepth2 [18] with ResNet50 encoder pre-trained on ImageNet [35] as the default depth estimator which is one of the most basic architectures in self-supervised MDE. For all experiments unless mentioned otherwise, we train the model for 50 epochs with a batch size of 8 and set $\tau_{\mathrm{mid}}$ to 20. We use Adam optimizer [28] with an initial learning rate of $5.0\times 10^{-5}$ decayed by half every 15 epochs. During training, we perform the following augmentations with 50% chance in random order: horizontal flips, brightness adjustment ($\pm$0.2), saturation adjustment ($\pm$0.2), contrast adjustment ($\pm$0.2), and hue jitter ($\pm$0.1). For quantitative evaluation, we follow [12] and compute the seven standard metrics: AbsRel, SqRel, RMSE, $\mathrm{RMSE_{log}}$, $\delta<1.25$, $\delta<1.25^{2}$, and $\delta<1.25^{3}$.
Table 1: Quantitative results on the KITTI Eigen test split [11] (640$\times$192 resolution). In the Supervision column, V, G, and CamH represent velocity, gravity, and camera height, respectively. The Scaling column represents whether to perform the median scaling [63]. We denote the best results in bold and the second best ones with an underline for each block. Note that we train VADepth [52] from scratch using ground-truth camera height (1.65m) provided by KITTI, unlike its original paper. FUMET outperforms weakly-supervised methods regardless of the network architectures and without the need for ground truth supervision. In the rescaled results, every model trained with FUMET achieves better results than its original one.
@@
Method Supervision Scaling AbsRel SqRel RMSE RMSElog $\delta\!\!<\!\!1.25$ $\delta\!\!<\!\!1.25^{2}$ $\delta\!\!<\!\!1.25^{3}$
@@
Figure 2: Qualitative comparison on KITTI. In error maps, the larger depth errors are represented in red, smaller ones are in blue. The model trained with our FUMET predicts more accurate depth maps compared to the weakly-supervised methods.
@@
Our FUMET is independent of network architectures so we experiment on several models with FUMET as well as Monodepth2: Lite-Mono [59] with ResNet18, HR-Depth [30] with ResNet18, and VADepth [52]. We train these models on the KITTI raw dataset [15], which contains 39,810 images for training and 4,424 for validation, and evaluate it with the KITTI Eigen test split [11] (697 images). For training and evaluation, all images are resized to 640$\times$192.
We show the qualitative results in Fig. 2. The quantitative results shown in Tab. 1 demonstrate that FUMET enables the various architectures to learn metric scale properly. Even when employing the simplest network architecture, i.e., Monodepth2, it outperforms weakly-supervised methods that rely on ground-truth scale labels for training which is not necessary for our method. Similar to VADepth [52], FUMET uses the camera height invariance as supervision though VADepth requires the ground-truth camera height and employs a complex network architecture. FUMET with Monodepth2 (row 6 in Tab. 1) achieves higher accuracy. This indicates that measuring camera height precisely is difficult and using unreliable ground-truth camera height causes low accuracy in the scale of estimated depth maps.
In the results showing the accuracy of the estimated depth map scaled with ground-truth scale factors via median scaling [63], it is evident that every model trained with FUMET attains superior results to the one trained without FUMET. This shows that our training framework not only enables the models to learn metric scale but also enhances the geometric accuracy. This improvement owes to the camera height loss $\mathcal{L}_{\mathrm{cam}}$ and $\mathcal{L}_{\mathrm{aux}}$, which makes the model respect road regions and object regions, respectively.
FUMET successfully enables a model to learn metric scale, even though the training dataset of KITTI contains many frames with no or few detected cars. We detail this in the supplementary material.
@@
To evaluate the ability of FUMET to leverage any dataset for training, we train and evaluate it on the Cityscapes dataset [9]. We follow [63, 53, 55, 49] and collect 69,731 images for training and evaluate on the 1,525 test images. In this dataset, there are multiple sets of camera intrinsics. Since this strongly affects the metric depth estimation [54], we first align the focal lengths of all images to 587.5 by resizing them, and crop them to 512$\times$192 resolution. For evaluation, we follow [5, 49] and further crop the estimated image into 416$\times$128. In order to align the number of iterations on KITTI, we train for 29 epochs with the same learning rate decreased by half every 8 epochs, and $\tau_{\mathrm{mid}}$ is set to 12.
Tab. 3 shows the quantitative results and FUMET outperforms weakly-supervised methods by wide margins. These methods assume that highly accurate sensor data is available and fail to train scale-aware MDE models on datasets consisting of unreliable sensor data such as Cityscapes. On the other hand, FUMET does not require any sensor measurement other than RGB video, so it can robustly leverage any driving videos for training.
@@
To demonstrate that FUMET can train a model on a mixed set of datasets to attain strong generalizability, we train with 235,341 images collected from the Argoverse2 [50], Lyft [24], A2D2 [16], and DDAD [19] dataset and evaluate it on the KITTI Eigen test split [11]. Both images for training and evaluation are resized and cropped to 832$\times$512 resolution. The training images are the sets of video sequences and each sequence is captured with its own unique camera setting. We individually optimize each camera height for each sequence during training. We detail the dataset composition and training settings in the supplementary material.
As shown in Tab. 3, the model trained with FUMET on the large-scale mixed dataset achieves comparable accuracy to the one trained on KITTI despite the domain gap and attains high generalizability. The model trained on both the mixed dataset and KITTI outperforms the model trained only on KITTI. This suggests that training on various images enhances accuracy which is the strength of fully self-supervised learning methods.
We also show that FUMET can leverage in-the-wild videos for training by using camera intrinsics recovered with SLAM, e.g., COLMAP [37] in the supplementary material.
Table 2: Quantitative results on Cityscapes [9]. We train all the models from scratch with the same supervisions as the Supervision column in Tab. 1. FUMET significantly outperforms weakly-supervised methods.
@@
FUMET 0.125 1.288 6.359 0.858
Table 3: FUMET can be trained on mixed datasets (Argoverse2, Lyft, A2D2, and DDAD) with various camera settings, and attains high generalizability. The results are evaluated on the KITTI Eigen split [11] (832$\times$512 resolution).
@@
To study the effect of the proposed components in FUMET, we conduct ablation studies on the training losses, balancing the loss weights, camera height optimization, and outlier filtering. The results are shown in Tab. 5. Thanks to the accuracy of LSP and the robustness of Silhouette Projector, the camera height loss makes a greater contribution to the depth accuracy than the auxiliary rough geometric loss. By introducing the camera height optimization, the model can learn the scale with more accurate and consistent supervision than leveraging the size prior independently for each frame (rows 1, 2, and 6 in Tab. 5), which leads to improved accuracy. Though balancing the loss weights (Eq. 14) with either $\mathcal{L}_{\mathrm{aux}}$ or $\mathcal{L}_{\mathrm{cam}}$ degrades the results due to the insufficient strength of the supervisory signals for metric scale, it contributes to improving accuracy when both $\mathcal{L}_{\mathrm{aux}}$ and $\mathcal{L}_{\mathrm{cam}}$ are employed.
Our proposed method can also be used with a pre-trained MDE model in an offline manner. We first train the original Monodepth2 [18] without FUMET, then predict depth maps of training images. With Silhouette Projector and LSP, we obtain per-frame scaled camera height and optimize it across frames with the median operation. We fix this value and use it as the camera height supervision for an untrained Monodepth2 on FUMET. We examine the difference between the proposed camera height optimized online and the fixed one derived offline when we use them as the pseudo-supervision. Note that we also fix $\lambda_{\mathrm{cam}}$ because the pseudo-label can be considered sufficiently accurate from the early stage of training. We also run another experiment in which we fix camera height supervision derived offline for the first 20 epochs, which is the same number of training epochs as the original Monodepth2, then unfix the camera height pseudo-label to fine-tune it via online optimization. As shown in Tab. 5, the model supervised with the fixed camera height computed offline achieves slightly better results than the one supervised with the camera height optimized online. This result indicates that our strategy of deriving scaled camera height is highly accurate and fixing it stabilizes the training. The model trained with the fine-tuned camera height achieves the highest accuracy, which shows that our FUMET can attain better accuracy with additional pre-training costs.